### v1

# Mixed Precision Iterative Refinement for Linear Inverse Problems

Abstract
This study investigates the iterative refinement method applied to the solution of linear discrete inverse problems by considering its application to the Tikhonov problem in mixed precision. Previous works on mixed precision iterative refinement methods for the solution of symmetric positive definite linear systems and least-squares problems have shown regularization to be a key requirement when computing low precision factorizations. For problems that are naturally severely ill-posed, we formulate the iterates of iterative refinement in mixed precision as a filtered solution using the preconditioned Landweber method with a Tikhonov-type preconditioner. Through numerical examples simulating various mixed precision choices, we showcase the filtering properties of the method and the achievement of comparable or superior accuracy compared to results computed in double precision as well as another approximate method.
keywords:
mixed precision, iterative refinement, Tikhonov regularization, preconditioned iterative methods, Landweber
{MSCcodes}
AMS Subject Classifications: 15A29 65F10 65F22 65F08
1 Introduction
We are interested in computing approximate solutions to linear least-squares problems of the form
(1) $\min_{x\in\mathbb{R}^{n}}\left\|Ax-b\right\|$
[…]
In applications the vector $b$ in (1) often represents measurements that are corrupted by error arising through noise contamination or instrumentation disturbances. From a numerical point of view this error may also include truncation error or approximation error. We denote this error by $e$ so that
$b=\hat{b}+e$
where $\hat{b}$ is the unknown error-free vector associated with $b$. Ideally, we would like to determine the solution of minimal Euclidean norm, $x^{\dagger}$, of the unavailable least-squares problem
(2) $x^{\dagger}=\arg\min_{x\in\mathbb{R}^{n}}\left\|Ax-\hat{b}\right\|.$
It’s well understood that because $b$ is contaminated by error and the singular values of $A$ cluster at the origin that the solution of (1) of minimal Euclidean norm is usually a poor approximation to $x^{\dagger}$. Regularization is a common strategy that replaces the problem (1) with a nearby problem that is less sensitive to the error in $b$.
In response to advancements in GPU computing architectures that natively support mixed precision computing, Higham and Pranesh in [18] consider the solution of symmetric positive definite (SPD) linear systems and least-squares problems in multiple precisions using iterative refinement (IR). They found that definiteness of SPD matrices could be lost when rounded to a precision with fewer representation bits, which in turn, could result in the failure of a Cholesky factorization when adapting GMRES-based IR strategies; see [2, 3]. It was determined that the definiteness of the matrix could be preserved by shifting the diagonal either before or after rounding - which may be viewed as a form of regularization. Our interests in this work are focused on applying IR to least-squares problems whose matrix $A$ is naturally severely ill-conditioned without any assumed underlying structure, to which Tikhonov regularization is often applied. Other works by Higham and others concerning the use of IR to solve linear systems using mixed precision include [12, 19, 23, 4].
In Tikhonov regularization, (1) is replaced by the penalized least-squares problem
(3) $\min_{x\in\mathbb{R}^{n}}\left\{\left\|Ax-b\right\|^{2}+\alpha^{2}\left\|Lx\right\|^{2}\right\},$
where $\alpha>0$ is a regularization parameter and $L\in\mathbb{R}^{s\times n}$ is a regularization matrix. We focus on what is termed standard Tikhonov, i.e. when $L=I$. When $L$ is chosen so that the null spaces of $A$ and $L$ trivially intersect then the solution of (3) may be expressed as
(4) $x^{(\alpha)}=\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}b,$
where the superscript T denotes the transpose operation. The regularization parameter $\alpha$ may be thought of as a ‘trade-off’ parameter which balances the sensitivity of the solution vector $x^{(\alpha)}$ to the error in $b$, as well as the closeness to the desired solution $x^{\dagger}$.
Given the unexpected necessity of regularization in structure leveraging algorithms for SPD linear systems in [18], we investigate the use of IR to solve the Tikhonov problem in mixed precision providing a natural extension to the aformentioned works. To better understand the regularized solutions of ill-posed problems computed using IR we derive a methodology to formulate the iterates applied to the Tikhonov problem as filtered solutions by writing them as a recursive relationship between the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous iterates. Additionally, we demonstrate in our numerical results that mixed precision IR on the Tikhonov problem gives comparable or superior accuracy against results computed in double precision as well as another benchmark which supports its use in modern applications that natively support mixed precision floating-point arithmetic.
An outline of this paper is as follows. In Section 2 we provide necessary background on topics including the Landweber iteration, IR, and filtering methods. Section 3 discusses preconditioned Landweber and its connection to the filtering analysis of IR on the Tikhonov problem in mixed precision, which is discussed in Section 4. Experimental results and concluding remarks are given in Sections 5 and 6, respectively.
2 Background
This section provides necessary background and notation for representing the regularized solution of a linear discrete ill-posed problem as a filtered solution. The Landweber and IR methods are also relevantly reviewed.
2.1 Filtered Solution of Linear Discrete Ill-posed Problems
Many inverse problems whose approximate solution is given by (1) have been well studied using the singular value decomposition (SVD) of $A=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices whose columns represent the left and right singular vectors of $A$, respectively. The diagonal matrix $\Sigma\in\mathbb{R}^{m\times n}$ contains the singular values on its main diagonal ordered in non-increasing fashion. Using the SVD, one can write the least-squares solution of (1) as a linear combination of the right singular vectors with the $j^{th}$ coefficient given by $u_{j}^{T}b/\sigma_{j}$. Since the singular values of $A$ cluster near numerical zero, the error $e$ contaminating $b$ is propagated into the computed solution for high index values $j$ corresponding to high frequency information [15, 14].
To prevent the propagation of error into the approximate solution of (1), the conditioning of the problem can be improved by employing filtering methods which can be formulated as a modified inverse solution [15, 22]. A filtered solution is of the form
(5) $x_{filt.}=\sum_{j=1}^{n}\phi_{j}\frac{u_{j}^{T}b}{\sigma_{j}}v_{j}$
where an intelligent selection of the filter factors $\phi_{j}\in\mathbb{R}$ can remove deleterious components of the approximate solution corresponding to high frequency information that is dominated by noise - which can be considered a regularized solution. To achieve an effective solution this corresponds to $\phi_{j}\approx 1$ for small $j$ and $\phi_{j}\approx 0$ for large $j$. Using the SVD of $A$, (4) may be written as a filtered solution with filter factors given by $\phi_{j}=\sigma_{j}^{2}/(\sigma_{j}^{2}+\alpha^{2})$ for $j=1,2,\dots,n$.
2.2 Landweber Iteration
A regularized solution of (1) can also be obtained by applying an iterative method and terminating the iterations before the error is propagated. The discrete Landweber iteration [21] is one such method whose $k^{th}$ iterate is given by
(6) $x^{(k)}=x^{(k-1)}+\zeta A^{T}\left(b-Ax^{(k-1)}\right)$
where $\zeta\in(0,2\sigma_{1}^{2})$ is a relaxation term with $\sigma_{1}$ the largest singular value of $A$. Without loss of generality, we choose $\zeta=1$ in this work since $A$ can be scaled appropriately. While robust, this classical variant of Landweber is well known to have slow convergence towards a useful solution which can be understood from viewing its $k^{th}$ solution as a filtering method [13]:
(7) $x^{(k)}=\sum_{j=1}^{n}\left(1-(1-\sigma_{j}^{2})^{k}\right)\frac{u_{j}^{T}b}{\sigma_{j}}v_{j}$
where we denote $\phi^{(k)}_{j}=\left(1-(1-\sigma_{j}^{2})^{k}\right)\in\mathbb{R}$ as the $j^{th}$ filter factor of the $k^{th}$ iterate.
For a scaled matrix $A$ whose largest singular value is one, only the first filter factor of the $k^{th}$ iteration, $\phi^{(k)}_{1}$ will be $1$, with subsequent filter factors of the same iteration rapidly decaying to numerical zero. It is only when $k$ grows that early filtering values, i.e. $\phi^{(k)}_{j}$ for small $j$, cluster near $1$. Because of this behavior, it can take many iterations to fully capture the dominant right singular vectors that constitute a meaningful approximate solution for inverse problems. One common way to overcome this difficulty is to employ a preconditioner; we comment more on this in Section 3.
2.3 Iterative Refinement on the Tikhonov Problem
Variants of fixed and mixed precision iterative refinement (IR) methods are not new - having been investigated consistently since Wilkinson’s first programmed version in 1948 [31]. A historical treatise of the progression of IR for linear problems may be found in [2, 16]. To compute the $k^{th}$ approximate solution, $x^{(k)}$, to the available linear system the $(k-1)^{th}$ residual $r^{(k-1)}=b-Ax^{(k-1)}$ is first computed. Then the correction equation $Ah=r^{(k-1)}$ is solved and the current solution is updated $x^{(k)}=x^{(k-1)}+h^{(k-1)}$. The process may be repeated as necessary until a termination criterion is met. When considering the application of IR to the Tikhonov problem given by
(8) $(A^{T}A+\alpha^{2}I)x^{(k)}=A^{T}b$
which was first discussed by Riley in [27], Golub in [11] noted that the IR procedure is equivalent to iterated Tikhonov regularization whose $(k+1)^{th}$ iterate is given by
(9) $x^{(k+1)}=x^{(k)}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k)}.$
The algorithm for IR on the Tikhonov problem is provided in Algorithm 1 for completeness.
Algorithm 1 Iterative refinement on the Tikhonov problem
1  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x^{(0)}\in\mathbb{R}^{n},\,\,\alpha>0$
2  Output: $x^{(k+1)}\in\mathbb{R}^{n}$
3  for $k=0,1,2,\dots$ do
4        $r^{(k)}=b-Ax^{(k)}$
5        $s^{(k)}=A^{T}r^{(k)}-\alpha^{2}x^{(k)}$
6        Solve $(A^{T}A+\alpha^{2}I)h^{(k)}=s^{(k)}$
7        $x^{(k+1)}=x^{(k)}+h^{(k)}$
8  end for
From Algorithm 1, the $k^{th}$ iterate of IR on the Tikhonov problem may be written recursively as
$x^{(k)}=x^{(k-1)}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(A^{T}A+\alpha^{2}I\right)^{-1}x^{(k-1)}.$
Here, each iterate involves the computation of a Tikhonov regularized linear system with $A$ and $r^{(k-1)}$. The $k^{th}$ filter factors take the form
(10) $\phi_{j}^{(k)}=1-\left(1-\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+\alpha^{2}}\right)^{k},\quad\quad j=1,2,\dots,n$
which are described in [11] and [20]. Differently than discussed in the aforementioned works, we show that the filter factors of IR on the Tikhonov problem can be written with respect to the $k^{th}$ iterate of iterated Tikhonov regularization (9). This framework lends itself to the filter factor analysis in mixed precision where instead of the $k^{th}$ iterate of (9) we utilize the $k^{th}$ iterate of preconditioned Landweber with a Tikhonov-type preconditioner, which we show in the next section can be written as a filtered solution.
3 Landweber with a Tikhonov-type Preconditioner
To consider the filtering of mixed precision IR on the Tikhonov problem we first consider the preconditioning of the Landweber iteration. Denote the right-preconditioned least-squares problem (1) by
(11) $\min_{x\in\mathbb{R}^{n}}\left\|AM^{-1}Mx-b\right\|$
with nonsingular preconditioner $M\in\mathbb{R}^{n\times n}$. By denoting $\hat{A}=AM^{-1}$ and $\hat{x}=Mx$ we may use Landweber (6) with $\zeta=1$ to solve (11), which may be simplified to
(12) $\hat{x}^{(k+1)}=x^{(k)}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax^{(k)}\right).$
We note that if $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$ then the preconditioned Landweber method is equivalent to iterated Tikhonov regularization (9). In this specific case, utilizing the SVD of $A$ we can represent $M^{T}M$ as follows
(13) $\begin{split}M^{T}M&=\left(V\Sigma^{T}U^{T}U\Sigma V^{T}+\alpha^{2}I\right)\\ &=V\left(\Sigma^{T}\Sigma+\alpha^{2}I\right)V^{T}\\ &=VD^{2}V^{T}\end{split}$
where $D^{2}=\left(\Sigma^{2}+\alpha^{2}\right)$ and $\Sigma^{2}=\Sigma^{T}\Sigma\in\mathbb{R}^{n\times n}$. Thus, one may consider the preconditioning matrix $M$ of (11) as $M=\left(\Sigma^{2}+\alpha^{2}\right)^{1/2}V^{T}=DV^{T}$: the product of a diagonal matrix with the right singular matrix of $A$.
Herein, we utilize the $M$ notation to help simplify our notation throughout this work and to consider the situation that $M^{T}M$ represents an approximate Tikhonov-type preconditioner to $A^{T}A+\alpha^{2}I$. Possible approximations we consider in Section 5 include that of (1) $M^{T}M$ as a low precision SVD approximation and (2) when the structure of $A$ allows for the approximation by a second matrix $C$ whose structure allows for fast and robust matrix-vector products. The former will be our focus when considering IR on the Tikhonov problem in mixed precision. The later case can arise naturally, for example, in image deblurring problems; see Section 5.
We clarify that there are various possible preconditioning strategies other than Tikhonov that could be used. It should be understood by our usage of the wording ‘preconditioned Landweber’ in this work that we only consider Tikhonov-type preconditioners. For further discussions and other preconditioner types considered for the solution of linear discrete ill-posed problems see [25, 1, 28].
To investigate the filter factors of mixed precision IR on the Tikhonov problem we first describe in Proposition 1 the filter factors of preconditioned Landweber with $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$. We again emphasize that the main use of the $M$ notation is to delineate between the possible change in structure of the preconditioner $M$ or its representation in a lower precision. Throughout, we will use a capital letter subscript, Q, to denote terms that come from a representation of $Q$, e.g. the SVD of $Q$. Additionally, we utilize subscripts $PL$ and $IR$ on vectors to denote solutions from preconditioned Landweber and IR, respectively.
Proposition 1.
(Filtered solution of preconditioned Landweber)
The $k^{th}$ iterative solution of (12) with preconditioner $M=D_{M}V_{M}^{T}$ as defined above may be written as
(14) $x^{(k)}_{PL}=\sum_{j=1}^{n}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{j}}v_{M,j}$
with the $j^{th}$ filter factor per $k^{th}$ iterative step given by
(15) $\psi^{(k)}_{j}=1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k},\quad j=1,2,\dots,n$
under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
Proof 3.1.
We begin by considering Landweber with $\zeta=1$ applied to our right-preconditioned least-squares problem (11) which may be represented by
$\hat{x}^{(k+1)}_{PL}=\hat{x}^{(k)}_{PL}+\hat{A}^{T}\left(b-\hat{A}\hat{x}^{(k)}\right)$
where we recall that $\hat{A}=AM^{-1}$ and $\hat{x}=Mx$. Using our definition of $M$ as $M=D_{M}V_{M}^{T}$ and the SVD of $A$, we may rewrite $\hat{A}$ as
$\hat{A}=U_{A}\Sigma_{A}V^{T}_{A}V_{M}D_{M}^{-1}.$
We may then write
$\hat{A}^{T}\hat{A}=D_{M}^{-1}V_{M}^{T}V_{A}\Sigma_{A}^{T}\Sigma_{A}V_{A}^{T}V_{M}D_{M}^{-1}$
so that
(16) $\begin{split}\hat{x}^{(k)}_{PL}&=\sum_{i=0}^{k-1}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b\\ &=\sum_{i=0}^{k-1}\left(I-D_{M}^{-1}V_{M}^{T}V_{A}\Sigma_{A}^{T}\Sigma_{A}V_{A}^{T}V_{M}D_{M}^{-1}\right)^{i}D_{M}^{-1}V_{M}^{T}V_{A}\Sigma_{A}^{T}U_{A}^{T}b.\end{split}$
Under our assumption, (16) simplifies to
$\hat{x}^{(k)}_{PL}=\sum_{i=0}^{k-1}\left(I-D_{M}^{-1}\Sigma_{A}^{T}\Sigma_{A}D_{M}^{-1}\right)^{i}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b.$
Denoting $\hat{D}^{k-1}=\sum_{i=0}^{k-1}\left(I-D_{M}^{-1}\Sigma_{A}^{T}\Sigma_{A}D_{M}^{-1}\right)^{i}$, we may write in terms of the $k^{th}$ iterative solution, $x^{(k)}$, by using that $x^{(k)}_{PL}=M^{-1}\hat{x}^{(k)}_{PL}$ so that
(17) $\begin{split}x^{(k)}_{PL}&=M^{-1}\hat{x}^{(k)}_{PL}\\ &=V_{M}D_{M}^{-1}\hat{D}^{k-1}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b\\ &=\sum_{j=1}^{n}\hat{d}_{j}^{k-1}\frac{\sigma_{j}}{d_{M,j}^{2}}\left(u_{A,j}^{T}b\right)v_{M,j}\end{split}$
where $\hat{d}_{j}^{k-1}\in\mathbb{R}$ denotes the $j^{th}$ diagonal entry of $\hat{D}^{k-1}$ and may be written as a geometric series
$\begin{split}\hat{d}_{j}^{k-1}&=\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}\right)^{i}\\ &=\frac{1-\left(1-\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}\right)^{k}}{\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}}.\end{split}$
With this we may write $x^{(k)}_{PL}$ as
$\begin{split}x^{(k)}_{PL}&=\sum_{j=1}^{n}\hat{d}_{j}^{k-1}\frac{\sigma_{j}}{d_{M,j}^{2}}\left(u_{A,j}^{T}b\right)v_{M,j}\\ &=\sum_{j=1}^{n}\left(1-\left(1-\frac{\sigma_{j}^{2}}{d_{M,j}^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{j}}v_{M,j}.\end{split}$
Taking the common denominator inside the term raised to the $k^{th}$ power and simplifying $d^{2}_{M,j}$ gives the filtered solution.
We mention several comments:
  1. 1.
    The orthogonality assumption between the right singular vectors of $A$ and $M$ is used in all results herein and may be (approximately) reasonable if the precision of $M$ relative to $A$ is close enough when $M^{T}M$ represents a low precision approximation to $A^{T}A+\alpha^{2}I$.
  2. 2.
    The solution of $x^{(k)}_{PL}$ is written using the basis vectors of $V_{M}$ instead of $V_{A}$ as in (7). Consider that we let $\hat{x}$ be the approximate solution of (11) which can be written as a filtered solution (5) using the SVD of $\hat{A}$. Then by definition of $M$, we can write $x=V_{M}D_{M}^{-1}\hat{x}$ showing that the basis choice of the solution method is independent of our orthogonality assumption.
  3. 3.
    For the remainder of this work we utilize the notation $\psi^{(k)}_{j}\in\mathbb{R}$ to denote the $j^{th}$ filter factor from the $k^{th}$ step of preconditioned Landweber and $\Psi^{(k)}\in\mathbb{R}^{n}$ to denote the vector containing the $n$ filter factors of the $k^{th}$ step.
4 Filtering Analysis of Iterative Refinement
In this section we derive the filter factors of IR on the Tikhonov problem using the filter factors from preconditioned Landweber. We first consider the simple case in Section 4.1 where a Tikhonov-type $M$ may be constructed from an approximate SVD of $A$ that could model a low-precision preconditioner. In Section 4.2 we extend our results to the determination of the filter factors in up to three precisions.
4.1 Reformulated Iterative Refinement on the Tikhonov problem
We begin by stating the $k^{th}$ solution of IR on the Tikhonov problem with preconditioner $M$ written recursively:
(18) $x^{(k)}_{IR}=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}.$
In Lemma 4.1, we show that the $k^{th}$ solution of IR on the Tikhonov problem may be written with respect to the $k^{th}$ solution of preconditioned Landweber given by (12).
Lemma 4.1.
(Reformulation of IR on the Tikhonov problem)
The $k^{th}$ solution of IR on the Tikhonov problem denoted by $x^{(k)}_{IR}$ may be written with respect to the $k^{th}$ solution of the preconditioned Landweber method (12) denoted by $x^{(k)}_{PL}$. Precisely, we have that
(19) $\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{k-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(k-1-i)}_{IR}.\end{split}$
Proof 4.2.
We proceed via induction defining that $x^{(0)}_{IR}=x^{(0)}_{PL}=0$.
Base case $(k=1)$:
Starting from the base relation of IR on the Tikhonov problem given by (18) we have that
$\begin{split}x^{(1)}_{IR}&=x^{(0)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(0)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(0)}_{IR}\\ &=\underbrace{x^{(0)}_{PL}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax^{(0)}_{PL}\right)}_{=x^{(1)}_{PL}}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(0)}_{IR}\\ &=x^{(1)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{0}x^{(0)}_{IR}\\ &=x^{(1)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{0}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(0-i)}_{IR}\end{split}$
completing the base case.
Inductive step:
Assume the inductive hypothesis is true for $k=n$ for $n>1$; that is
$\begin{split}x^{(n)}_{IR}&=x^{(n-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(n-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n-1)}_{IR}\\ &=x^{(n)}_{PL}-\underbrace{\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{n-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(n-1-i)}_{IR}}_{=\Gamma}.\end{split}$
We proceed as follows
$\begin{split}&x^{(n+1)}_{IR}=x^{(n)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax^{(n)}_{IR}\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-A\left[x^{(n)}_{PL}-\Gamma\right]\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n)}_{IR}+\underbrace{\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax_{PL}^{(n)}\right)}_{=h^{(n)}}-\left(M^{T}M\right)^{-1}A^{T}A\Gamma-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n)}_{PL}-\Gamma+h^{(n)}-\left(M^{T}M\right)^{-1}A^{T}A\Gamma-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n+1)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{0}x^{(n)}_{IR}\\ &\hskip 14.22636pt-\alpha^{2}\left(M^{T}M\right)^{-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)\sum_{i=0}^{n-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(n-1-i)}_{IR}\\ &=x^{(n+1)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{n}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(n-i)}_{IR}.\end{split}$
Since both the base case and the inductive step have been shown, by mathematical induction the relation (19) holds for every natural number $k$.
With the result of Lemma 4.1, we may observe that the $k^{th}$ iterate of IR on the Tikhonov problem may be computed using the first $k$ solutions of (14). As such, it is possible to recursively determine the filter factors for $x^{(k)}_{IR}$. Consider the first iteration of IR using the result from Lemma 4.1:
$\begin{split}x^{(1)}_{IR}&=x^{(1)}_{PL}-\alpha^{2}(M^{T}M)^{-1}\sum_{i=0}^{0}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(0-i)}_{IR}\\ &=x^{(1)}_{PL}=\sum_{j=1}^{n}\underbrace{\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)}_{=\phi_{j}^{(1)}}\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}.\end{split}$
Here and throughout, we differentiate the filter factors of IR from preconditioned Landweber by defining that the filter factors of $x^{(k)}_{IR}$ are given by $\Phi_{IR}^{(k)}\in\mathbb{R}^{n}$ whose $j^{th}$ term we denoted by $\phi_{j}^{(k)}$. To determine the filter factors of $x^{(2)}_{IR}$ we proceed as follows using that $x^{(1)}_{IR}=x^{(1)}_{PL}$
(20) $\begin{split}x^{(2)}_{IR}&=x^{(2)}_{PL}-\alpha^{2}(M^{T}M)^{-1}\sum_{i=0}^{1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(1-i)}_{IR}\\ &=x^{(2)}_{PL}-\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{0}x^{(1)}_{PL}.\end{split}$
Before proceeding, we consider the general situation that will arise when expanding the $k^{th}$ solution of IR with respect to the solutions of preconditioned Landweber. That is, we consider the simplification of
(21) $\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(q)}_{PL},\quad\begin{cases}i=0,1,\dots,k-1\\ q=0,1,\dots,k-1\end{cases}.$
We first note that $\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}$ may be decomposed using the definition of the preconditioner $M$, the SVD of $A$, and the orthgonality assumption between the right singular vectors of $A$ and of $M$:
$\begin{split}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}&=\left(I-V_{A}\Sigma_{A}^{2}V_{A}^{T}V_{M}D_{M}^{-1}V_{M}^{T}\right)^{i}\\ &=V_{A}\left(I-\tilde{D}_{M}\right)^{i}V_{M}^{T}\end{split}$
where $\tilde{D}_{M}=\Sigma_{A}^{2}D_{M}^{-1}=\text{diag}\left(\sigma_{A,1}^{2}/(\sigma_{M,1}^{2}+\alpha^{2}),\dots,\sigma_{A,n}^{2}/(\sigma_{M,n}^{2}+\alpha^{2})\right)$. With this, we may rewrite (21) following the relation given by (17)
$\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(q)}_{PL}$
$=\alpha^{2}V_{M}D_{M}^{-1}(I-\tilde{D}_{M})^{i}D_{M}^{-1}\hat{D}^{k-1}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b$
$=\sum_{j=1}^{n}\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}.$
Returning to (20), we may rewrite $x^{(2)}_{IR}$ as follows
$\begin{split}x^{(2)}_{IR}&=x^{(2)}_{PL}-\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{0}x^{(1)}_{PL}\\ &=\sum_{j=1}^{n}\phi^{(2)}_{j}\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}\end{split}$
where
$\phi^{(2)}_{j}=\psi^{(2)}_{j}-\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{0}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)$
for $j=1,2,\dots,n$. Here, we note that $\Phi^{(2)}$ is the difference between the filter factors of $x^{(2)}_{PL}$ and a scaling of $\Phi^{(1)}$. Continuing this recursive process, we may arrive at the following general result.
Theorem 4.3.
(IR on Tikhonov problem with filter factors)
The $k^{th}$ solution of IR on the Tikhonov problem, $x^{(k)}_{IR}$, may be written as
$x^{(k)}_{IR}=\sum_{j=1}^{n}\left(\psi^{(k)}_{j}-\left(\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-1-i)}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}$
with the $j^{th}$ filter factor per $k^{th}$ step given by
$\phi_{j}^{(k)}=\psi_{j}^{(k)}-\left(\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-1-i)},\quad j=1,2,\dots,n$
under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
We comment that the filter factors of IR on the Tikhonov problem computed in exact arithmetic are exactly that of Tikhonov regularization. This is stated in Corollary 4.4 below without proof as the argument follows from a straight forward mathematical induction argument.
Corollary 4.4.
(Filter factors of IR on the Tikhonov problem in exact arithmetic)
In exact arithmetic the filter factors of IR on the Tikhonov problem coincide with the filter factors of standard Tikhonov regularization (4) for all $k=0,1,2,\dots$ given by
$\phi^{(k)}_{j}=\frac{\sigma^{2}_{j}}{\sigma^{2}_{j}+\alpha^{2}},\quad\text{for}\,\,\,j=1,2,\dots,n.$
4.2 Mixed precision IR on the Tikhonov problem
We now turn to the determination of the filter factors of IR when computing in multiple precisions. We extend the result of Theorem 4.3 to the case when the $k^{th}$ iterate and its corresponding filter factors are to be computed in up to three precisions. When discussing computations in mixed precision we match the convention of recent works, e.g., [12, 16], by using the notation
$\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
to represent the three precisions used. Here, Pr3 denotes the highest precision that is used in residual updates, Pr2 denotes the ‘working precision’ used for linear systems solves, and Pr1 denotes the precision used to represent the preconditioner $M$. In Section 5 we define the precisions we consider for our numerical experiments.
Before extending Theorem 4.3, we present the generalization of Proposition 1 to the mixed precision case. We provide in Algorithm 2 the steps with their associated precisions for computing preconditioned Landweber. In Proposition 2 we highlight the parts of the filtered solution that are computed in Pr1, Pr2, and Pr3. The framework for the computations of the filter factors is summarized in Algorithm 3.
Algorithm 2 Mixed precision preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x^{(0)}=0\in\mathbb{R}^{n},\,\,\alpha>0$
3  Output: $x^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        $s^{(k)}=A^{T}\left(b-Ax^{(k)}\right)$$[\text{in }\text{Pr}_{3}]$
7        Solve $(M^{T}M)h^{(k)}=s^{(k)}$ $[\text{in }\text{Pr}_{2}]$
8        $x^{(k+1)}=x^{(k)}+h^{(k)}$ $[\text{in }\text{Pr}_{2}]\,\,$
9  end for
Proposition 2.
(Mixed precision preconditioned Landweber with filter factors)
The $k^{th}$ solution, $x^{(k)}_{PL}$, of (12) with preconditioner $M=D_{M}V_{M}^{T}$ as defined in Section 2 in precisions Pr1, Pr2, and Pr3 may be written as
(22) $x^{(k)}_{PL}=\sum_{j=1}^{n}\left(\psi_{j}^{(k-1)}+\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k-1}\right)\frac{u^{T}_{A,j}b}{\sigma_{A,j}}v_{M,j}$
with filter factors per $k^{th}$ step given by
(23) $\psi_{j}^{(k)}=\psi_{j}^{(k-1)}+\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k-1}\quad\text{for}\,\,\,j=1,2,\dots,n$
under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
Proof 4.5.
From (12) we have that
$x^{(k)}_{PL}=x^{(k-1)}_{PL}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}.$
To write the $k^{th}$ iterate as a filtered solution we rewrite $A^{T}r^{(k-1)}$ while retaining that it be computed in Pr3 from Algorithm 2. Using (6) and (16) with $\zeta=1$ we may write
$\begin{split}A^{T}r^{(k-1)}&=\left(M^{T}M\right)\left(x^{(k)}_{PL}-x^{(k-1)}_{PL}\right)\\ &=\left(M^{T}M\right)\left[M^{-1}\left(\hat{x}_{PL}^{k}-\hat{x}_{PL}^{k-1}\right)\right]\\ &=\left(M^{T}M\right)\left[M^{-1}\left(\sum_{i=0}^{k-1}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b-\sum_{i=0}^{k-2}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b\right)\right]\\ &=\left(M^{T}M\right)M^{-1}\left(I-\hat{A}^{T}\hat{A}\right)^{k-1}\hat{A}^{T}b.\end{split}$
Using the definition of $M$, $\hat{A}$, and the SVD of $A$ we can rewrite preconditioned Landweber using our orthogonality assumption as follows
$\begin{split}x^{(k)}_{PL}&=x^{(k-1)}_{PL}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}\\ &=x^{(k-1)}_{PL}+V_{M}D_{M}^{-2}D_{M}\left(I-D_{M}^{-1}\Sigma^{2}_{A}D_{M}^{-1}\right)^{k-1}D_{M}^{-1}\Sigma_{A}^{T}U_{A}^{T}b.\end{split}$
By defining the $j^{th}$ component of $\left(I-D_{M}^{-1}\Sigma^{2}_{A}D_{M}^{-1}\right)^{k-1}$ by $\tilde{d}_{j}^{k-1}=(1-\sigma_{A,j}^{2}/\sigma^{2}_{M,j}+\alpha^{2})^{k-1}$ which come from the computation of $A^{T}r^{(k-1)}$, we may write
$x^{(k)}_{PL}=x^{(k-1)}_{PL}+\sum_{j=1}^{n}\left(\left(d^{2}_{M,j}\right)^{-1}\sigma^{2}_{A,j}\tilde{d}^{k-1}_{j}\right)\frac{u^{T}_{A,j}b}{\sigma_{A,j}}v_{M,j}.$
Here, $d_{M,j}^{2}=(\sigma^{2}_{M,j}+\alpha^{2})$ represents the components from the preconditioner computed in Pr1 and whose component-wise action is computed in Pr2. Simplifying, we have
$x^{(k)}_{PL}=\sum_{j=1}^{n}\left(\psi_{j}^{(k-1)}+\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k-1}\right)\frac{u^{T}_{A,j}b}{\sigma_{A,j}}v_{M,j}$
where the update to $\psi_{j}^{(k-1)}$ is also computed in Pr2.
Algorithm 3 Mixed precision filter factors of preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\Sigma^{2}_{M}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
3  Output: $\Psi^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        for $j=1,2,\dots,n$ do
7              $q_{j}^{(k)}=\sigma_{A,j}^{2}\left(1-\sigma^{2}_{A,j}\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}\right)^{k}$ $[\text{in }\text{Pr}_{3}]$
8              $g_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}q_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]$
9              $\psi_{j}^{(k+1)}=\psi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]\,\,$
10        end for
11  end for
With Proposition 2 we now consider the reformulation of IR on the Tikhonov problem in mixed precision so as to determine its corresponding filter factors. We provide the mixed precision variant of Algorithm 1 in Algorithm 4 for reference. Recall that the $k^{th}$ solution of (18) can be written as a recursive relation involving the $k^{th}$ iterate of preconditoned Landweber and the previous $k-1$ iterates of IR on the Tikhonov problem. In Lemma 4.6 we reformulate (18) to write the $k^{th}$ iterate of the IR problem as a filtered solution whose constituent parts are computed in precisions that mimic their partnered forms in Algorithm 4.
Algorithm 4 Mixed precision iterative refinement on the Tikhonov problem
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,x^{(0)}\in\mathbb{R}^{n}$
3  Output: $x^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $M^{T}M=\left(A^{T}A+\alpha^{2}I\right);$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        $r^{(k)}=b-Ax^{(k)};$ $[\text{in }\text{Pr}_{3}]$
7        $s^{(k)}=A^{T}r^{(k)}-\alpha^{2}x^{(k)};$ $[\text{in }\text{Pr}_{3}]$
8        Solve $\left(M^{T}M\right)h^{(k)}=s^{(k)};$ $[\text{in }\text{Pr}_{2}]$
9        $x^{(k+1)}=x^{(k)}+h^{(k)};$ $[\text{in }\text{Pr}_{2}]\,\,$
10  end for
Lemma 4.6.
(Reformulation of IR on the Tikhonov problem for mixed precision)
The $k^{th}$ solution of IR on the Tikhonov problem, $x_{IR}^{(k)}$, in precisions Pr1, Pr2, and Pr3 may be written with respect to the $k^{th}$ solution of mixed precision preconditioned Landweber computed by Proposition 2. Precisely,
(24) $\begin{split}&x^{(k)}_{IR}=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}\left[\left(M^{T}M\right)\left[x^{(k)}_{PL}-x^{(k-1)}_{PL}\right]+\Gamma-\alpha^{2}x_{IR}^{(k-1)}\right]\end{split}$
where $\Gamma=A^{T}A\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{k-2}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x_{IR}^{(k-2-i)}$.
Proof 4.7.
We begin by considering the IR expansion about its $(k-1)^{th}$ residual
(25) $x^{(k)}_{IR}=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax_{IR}^{(k-1)}\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}.$
Using Lemma 4.1 we may rewrite the residual of $x^{(k-1)}_{IR}$ in (25) instead with respect to $x^{(k-1)}_{PL}$
$\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax_{IR}^{(k-1)}\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r_{PL}^{(k-1)}+\left(M^{T}M\right)^{-1}\Gamma-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}\left[A^{T}r_{PL}^{(k-1)}+\Gamma-\alpha^{2}x^{(k-1)}_{IR}\right]\end{split}$
where $r^{(k-1)}_{PL}$ denotes the $(k-1)^{th}$ residual of preconditioned Landweber solution given by (22). To compute $x^{(k)}_{IR}$ with respect to its own prior solutions and $x_{PL}^{(q)}$ for $q=0,1,\dots,k$, we rewrite $A^{T}r_{PL}^{(k-1)}$ using (12) as follows
$x_{PL}^{(k)}=x_{PL}^{k-1}+\left(M^{T}M\right)^{-1}A^{T}r_{PL}^{(k-1)}\\ \iff\\ A^{T}r_{PL}^{(k-1)}=\left(M^{T}M\right)\left[x_{PL}^{(k)}-x_{PL}^{(k-1)}\right].$
Here, we point out that by construction $M^{T}M$ is invertible. With this substitution the result follows.
We now describe the computation of the filter factors of IR on the Tikhonov problem in precisions Pr1, Pr2, and Pr3. The result is provided without proof as the derivations follow in a similar manner to Theorem 4.3 when starting with the result of Lemma 4.1. Here, we adhere to the same convention outlined in Proposition 2 regarding the computation of residuals, linear systems solves, and the preconditioner in appropriate precisions. Algorithm 5 summarizes the framework for computing the filter factors.
Theorem 4.8.
(Mixed precision IR on the Tikhonov problem with filter factors)
The $k^{th}$ solution of IR on the Tikhonov problem, $x^{(k)}_{IR}$, in precisions Pr1, Pr2, and Pr3 may be written as
$x^{(k)}_{IR}=\sum_{j=1}^{n}\phi_{j}^{(k)}\frac{u^{T}_{A,j}b}{\sigma_{A,j}}v_{M,j}$
with the $j^{th}$ filter factor per $k^{th}$ iterative step given by
$\begin{split}\phi_{j}^{(k)}=\phi_{j}^{(k-1)}+\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}&\Biggl{[}\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\left[\psi_{j}^{(k)}-\psi_{j}^{(k-1)}\right]-\alpha^{2}\phi^{(k-1)}_{j}\\ &+\left(\frac{\alpha^{2}\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-2}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-2-i)}\Biggr{]}\end{split}$
for $j=1,2,\dots,n$ under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
Algorithm 5 Mixed precision filter factors of iterative refinement on the Tikhonov problem
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Phi^{(0)}=0\in\mathbb{R}^{n},\,\,\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\Sigma^{2}_{M}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
3  Output: $\Phi^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        Compute $\Psi^{(k+1)}$ using Algorithm 3 $[\text{in }\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3}]$
7        for $j=1,2,\dots,n$ do
8              $q_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\left[\psi_{j}^{(k+1)}-\psi_{j}^{(k)}\right]-\alpha^{2}\phi^{(k-1)}_{j}$ $[\text{in }\text{Pr}_{3}]$
9              $q_{j}^{(k)}=q_{j}^{(k)}+\left(\frac{\alpha^{2}\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-2}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-2-i)}$ $[\text{in }\text{Pr}_{3}]$
10              $g_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}q_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]$
11              $\phi_{j}^{(k+1)}=\phi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]$
12        end for
13  end for
5 Numerical Results and Preliminaries
In this section we provide numerical experiments to illustrate the effectiveness of mixed precision IR (MP-IR) on the Tikhonov problem and provide an exposition of the filter factor results discussed in Section 4. The section is organized as follows: Sections 5.1 and 5.2 provide the background for the determination of what we will refer to as effective filter factors of IR on the Tikhonov problem and an overview of the simulation of low precision computations in MATLAB, respectively. In Section 5.3 we discuss our numerical results and their associated preliminaries.
5.1 Effective filter factors
To assess the experimental veracity of our filter factor results from Section 4 for IR on the Tikhonov problem we compare against experimentally derived effective filter factors. Using (5), the effective filter factors $\omega^{(k)}_{j}$ of an approximate solution at the $k^{th}$ iteration may be computed as follows
(26) $\omega^{(k)}_{j}=\frac{v_{M,j}^{T}x^{(k)}_{IR}}{u^{T}_{A,j}b}\sigma_{A,j},\quad j=1,2,\dots,n$
where we denote the vector containing the filter factors of the $k^{th}$ iterate by $\Omega^{(k)}\in\mathbb{R}^{n}$. We highlight that (26) uses the right singular vectors of the preconditioner $M$ as a basis of the $k^{th}$ iterate which we commented on in our remarks following Proposition 1. In Section 4 we showed that this also held true in Theorems 4.3 and 4.8.
5.2 Low precision simulation in MATLAB
To ascertain the effectiveness of mixed precision IR on the Tikhonov problem and the experimental validity of the results presented in Section 4 we utilize the software package chop introduced by Higham and Pranesh in [17]. The chop function simulates lower precision arithmetic by rounding array entries given in a MATLAB native precision (e.g. single or double) to a target precision which is stored in a higher precision with non-utilized representation bits set to zero. Chen et al. in [6] considered the application of chop for the solution of structured inverse problems as well as associated numerical considerations of utilizing chop.
Various target precisions are supported by the chop software including user-customizable ones. The precisions we utilize in our numerical experiments in Section 5.3 and their associated shorthand notations are provided in Table 1. We comment that we did not experience any significant experimental differences when using precisions that did not support subnormal numbers, e.g. Google’s bfloat16111https://en.wikipedia.org/wiki/Bfloat16__\__floating-point__\__format, compared to those due to the IEEE standard [^2] https://en.wikipedia.org/wiki/IEEE__\__754##\##2019 during our numerical investigations and therefore do not include them in our results.
Table 1: Precisions considered in the numerical results: shorthand given as an integer, bits given for the mantissa and exponent, and whether precision is IEEE standard or not.
Pr Shorthand Name Exponent Bits Mantissa Bits IEEE Std.
$1$ fp64 $11$ $52$ Yes
$2$ fp32 $8$ $23$ Yes
$3$ fp16 $5$ $10$ Yes
$4$ fp8 $4$ $3$ No (custom)
The reported numerical results of this work were carried out in MATLAB R2022b 64-bit on a MacBook Pro laptop running MacOS Ventura with an Apple M2 Pro processor with @3.49 GHz and 16 GB of RAM. The computations other than those utilizing chop were carried out with about 15 significant decimal digits.
5.3 Examples and Results
For our first example we consider the 1D signal restoration problem Spectra whose matrix models a symmetric Gaussian blur and $x$ is a simulated X-ray spectrum [30]. The $a_{i,j}$ entries of $A\in\mathbb{R}^{64\times 64}$ are given by
$a_{i,j}=\frac{1}{\eta\sqrt{2\pi}}\exp\left(-\frac{(i-j)^{2}}{2\eta^{2}}\right),$
with $\eta=2$ which results in a Toeplitz matrix.
To realistically simulate the inverse problem, noise was added to the true right-hand side, $\hat{b}$, by forming the vector $e$ with normally distributed random entries with mean zero so that $b=Ax+e$; the vector $e$ is scaled so as to correspond to a specific noise level given by $\mu=100\left(\|e\|/\|\hat{b}\|\right)$. We will refer to $\mu$ as the noise level. The condition number of the matrix $A$ as determined by the MATLAB function cond() is $\approx 10^{9}$; it can also be easily verified that the singular values of this matrix decay without a significant gap.
We compare the solution quality of MP-IR to what we will term the Approximated Iterative Refinement (AIR) method inspired by the work of Donatelli and Hanke in [8]. Here, we let the preconditioner be $M^{T}M=\left(C^{\ast}C+\alpha^{2}I\right)$ where $C$ represents a circulant approximation to $A$. From this structure, the preconditioner is normal and admits a spectral decomposition that can be computed efficiently. We approximate $C\approx A$ using the scheme devised by Chan in [5] which minimizes the Frobenius norm of the difference between $A$ and $C$ amongst a family of circulants. This approximation scheme was analyzed by Strela and Tyrtyshnikov [29] and found to be a good choice from the viewpoint of eigenvalue clustering when the vector to be reconstructed takes on many zero values.
(Spectra - filter factors in fp64) - We begin our filter factor analysis by evaluating the filtering behavior of MP-IR on the Spectra problem contaminated by $1\%$ noise in fp64. For this experiment we chose the regularization parameter $\alpha^{2}=1e-2$; in the restoration part of this example we expand on this choice. Figure 1 displays the filter factors from Theorem 4.3, Theorem 4.8, and the effective filter factors compared to those of Tikhonov when considering all computations in fp64 after $1$ iteration. We comment that there is no visual discernible difference amongst the plots thereby experimentally verifying that: (1) the filter factors of Theorem 4.3 and Theorem 4.8 coincide when all floating point computations are done in fp64, i.e. when $(\text{Pr}_{1},\,\text{Pr}_{2},\,\text{Pr}_{3})=(1,1,1)$ and (2) that all three sets of filter factors match those of Tikhonov. Table 2 (discussed below) numerically confirms the agreement between Theorem 4.8 and the effective filter factors in fp64 working precision.
Figure 1: Filter factors of the Spectra problem with $1\%$ noise compared against those of Tikhonov with $\alpha^{2}=1e-2$ after $1$ iteration using (a) the effective filter factors, (b) Theorem 4.3, and (c) Theorem 4.8.
[…]
Figure 2: MP-IR filter factors (bottom row) determined by Theorem 4.8 with $(\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3})=(3,2,1)$ for $1\%$ noise compared to their entry wise computations of $|\Phi^{(k)}-\Omega^{(k)}|$ (top row) at iteration numbers (a) 1, (b) 5, and (c) 10.
Table 2: Spectra example: summary statistics of the absolute difference between the effective filter factors and those computed by Theorem 4.8 given by $|\Phi^{(k)}-\Omega^{(k)}|$ for various precision combinations with a noise level of $1\%$ and $\alpha^{2}=$1e-2.
(Pr1, Pr2, Pr3) Iter. Summary Stats. of $|\Phi^{(k)}-\Omega^{(k)}|$
Mean Error Min Error Max Error SD Error
(1,1,1) $1$ $-141.5e\text{-}$ 14 $-205.5e\text{-}$ 20 $-142.4e\text{-}$ 14 $-143.8e\text{-}$ 14
$5$ $-154.9e\text{-}$ 15 00 $-132.0e\text{-}$ 13 $-142.5e\text{-}$ 14
$10$ $-154.7e\text{-}$ 15 00 $-131.9e\text{-}$ 13 $-142.4e\text{-}$ 14
(2,1,1) $1$ $-61.5e\text{-}$ 6 $-148.8e\text{-}$ 14 $-54.7e\text{-}$ 5 $-66.4e\text{-}$ 6
$5$ $-84.1e\text{-}$ 8 $-141.7e\text{-}$ 14 $-77.2e\text{-}$ 7 $-71.2e\text{-}$ 7
$10$ $-84.1e\text{-}$ 8 $-141.7e\text{-}$ 14 $-77.2e\text{-}$ 7 $-71.2e\text{-}$ 7
(2,2,1) $1$ $-61.5e\text{-}$ 6 $-142.5e\text{-}$ 14 $-54.8e\text{-}$ 5 $-66.4e\text{-}$ 6
$5$ $-85.2e\text{-}$ 8 $-157.7e\text{-}$ 15 $-76.7e\text{-}$ 7 $-71.3e\text{-}$ 7
$10$ $-85.6e\text{-}$ 8 $-157.7e\text{-}$ 15 $-76.7e\text{-}$ 7 $-71.4e\text{-}$ 7
(3,2,1) $1$ $-21.2e\text{-}$ 2 $-95.0e\text{-}$ 9 $-12.9e\text{-}$ 1 $-24.0e\text{-}$ 2
$5$ $-46.5e\text{-}$ 4 $-121.1e\text{-}$ 12 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
$10$ $-46.5e\text{-}$ 4 $-121.1e\text{-}$ 12 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
(3,2,2) $1$ $-21.2e\text{-}$ 2 $-95.1e\text{-}$ 9 $-12.9e\text{-}$ 1 $-24.0e\text{-}$ 2
$5$ $-46.5e\text{-}$ 4 $-132.4e\text{-}$ 13 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
$10$ $-46.5e\text{-}$ 4 $-121.0e\text{-}$ 12 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
(3,3,2) $1$ $-21.2e\text{-}$ 2 $-95.0e\text{-}$ 9 $-12.9e\text{-}$ 1 $-24.1e\text{-}$ 2
$5$ $-46.6e\text{-}$ 4 $-105.0e\text{-}$ 10 $-22.7e\text{-}$ 2 $-33.3e\text{-}$ 3
$10$ $-46.8e\text{-}$ 4 $-104.4e\text{-}$ 10 $-22.7e\text{-}$ 2 $-33.3e\text{-}$ 3
(4,3,2) $1$ $01.3e$ 0 $-61.0e\text{-}$ 6 $13.6e$ 1 $05.0e$ 0
$5$ $-23.7e\text{-}$ 2 $-84.0e\text{-}$ 8 $-18.5e\text{-}$ 1 $-11.3e\text{-}$ 1
$10$ $-23.5e\text{-}$ 2 $-83.1e\text{-}$ 8 $-19.9e\text{-}$ 1 $-11.4e\text{-}$ 1
(4,3,3) $1$ $01.3e$ 0 $-79.6e\text{-}$ 7 $13.6e$ 1 $05.0e$ 0
$5$ $-23.8e\text{-}$ 2 $-84.0e\text{-}$ 8 $-18.9e\text{-}$ 1 $-11.4e\text{-}$ 1
$10$ $-23.6e\text{-}$ 2 $-83.2e\text{-}$ 8 $01.0e$ 0 $-11.4e\text{-}$ 1
(4,4,3) $1$ $01.4e$ 0 $-61.0e\text{-}$ 6 $13.6e$ 1 $05.0e$ 0
$5$ $-11.0e\text{-}$ 1 $-82.8e\text{-}$ 8 $04.4e$ 0 $-15.5e\text{-}$ 1
$10$ $-24.8e\text{-}$ 2 $-81.6e\text{-}$ 8 $01.3e$ 0 $-11.7e\text{-}$ 1
(4,4,4) $1$ $02.0e$ 0 $-78.6e\text{-}$ 7 $14.7e$ 1 $07.4e$ 0
$5$ $-11.3e\text{-}$ 1 $-81.0e\text{-}$ 8 $04.2e$ 0 $-15.3e\text{-}$ 1
$10$ $-11.4e\text{-}$ 1 $-91.5e\text{-}$ 9 $02.6e$ 0 $-13.9e\text{-}$ 1
(Spectra - reconstruction) - We now compare MP-IR versus AIR on the Spectra problem. As our interests have focused on the filtering properties of MP-IR on the Tikhonov problem, we did not investigate methodologies for determining regularization parameters nor did we consider any stopping criterion. Both topics deserve their own dedicated studies in the context of mixed precision computing for inverse problems. In our numerical investigation we chose 3 magnitudes of regularization parameters: 1e-1, 1e-2, and 1e-3 as well as two noise levels: $0.5\%$ and $3\%$ to consider the broad behavior of both methods. For method comparison, we compute the relative reconstructive error (RRE) defined by
$\text{RRE}\left(x^{(k)}\right)=\frac{\left\|x^{(k)}-x\right\|}{\|x\|}$
where $x$ and $x^{(k)}$ denote the true solution and $k^{th}$ approximate solution determined by MP-IR or AIR, respectively.
Table 3 provides the RREs at selected iterations and precisions for MP-IR as well as AIR as a function of $\alpha^{2}$ and the noise level of the problem. For succinctness in the table, we did not provide the results for $\alpha^{2}=$1e-1 as AIR and all MP-IR precision variants did about the same in terms of minimum and final RREs amongst the $10$ iterations considered for both noise levels. Specifically, the RREs for both methods were $\approx 0.62$.
When $\alpha^{2}$ was decreased by an order of magnitude we observed that AIR attained larger errors than all MP-IR variants and was not stable as evident by a large final RRE value at iteration $10$. This general behavior was also observed when $\alpha^{2}=$1e-3 indicating that MP-IR would be a superior method for these choices of $\alpha^{2}$. We also observed that as the 3 precision choices for MP-IR were changed away from fp64 that it would take more iterations to reach the minimum RRE value. However, the relative errors still remained stable compared to AIR for the same number of iterations.
Table 3: Spectra example: relative errors for MP-IR and AIR methods as a function of the regularization parameter and noise level. The minimum error and its corresponding iteration number in parentheses as well as the final error for the maximum number of iterations allowed (10) is provided. MP-IR errors are provided for various mixed precision combinations.
$\alpha^{2}$ Method (Pr1, Pr2, Pr3) Noise Level
$0.5\%$ $3\%$
Min. RRE (It.) Fin. RRE Min. RRE (It.) Fin. RRE
1e-2 AIR $0.504\,\,(1)$ $46.9e$ 4 $0.517\,\,(1)$ $54.2e$ 5
MP-IR (1,1,1) $0.497\,\,(7)$ $0.497$ $0.500\,\,(1)$ $0.500$
(2,1,1) $0.497\,\,(1)$ $0.497$ $0.500\,\,(3)$ $0.500$
(2,2,1) $0.497\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(3,2,1) $0.496\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(3,2,2) $0.496\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(3,3,2) $0.496\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(4,3,2) $0.497\,\,(3)$ $0.497$ $0.500\,\,(2)$ $0.500$
(4,3,3) $0.496\,\,(3)$ $0.496$ $0.500\,\,(6)$ $0.500$
(4,4,3) $0.497\,\,(2)$ $0.497$ $0.501\,\,(7)$ $0.501$
(4,4,4) $0.496\,\,(10)$ $0.496$ $0.501\,\,(4)$ $0.501$
1e-3 AIR $0.406\,\,(1)$ $127.4e$ 12 $0.998\,\,(1)$ $134.4e$ 13
MP-IR (1,1,1) $0.347\,\,(2)$ $0.347$ $0.373\,\,(1)$ $0.373$
(2,1,1) $0.347\,\,(10)$ $0.347$ $0.373\,\,(3)$ $0.373$
(2,2,1) $0.347\,\,(2)$ $0.347$ $0.373\,\,(2)$ $0.373$
(3,2,1) $0.347\,\,(3)$ $0.347$ $0.373\,\,(3)$ $0.373$
(3,2,2) $0.347\,\,(3)$ $0.347$ $0.373\,\,(3)$ $0.373$
(3,3,2) $0.347\,\,(7)$ $0.347$ $0.373\,\,(8)$ $0.373$
(4,3,2) $0.346\,\,(2)$ $0.347$ $0.373\,\,(6)$ $0.373$
(4,3,3) $0.346\,\,(5)$ $0.347$ $0.373\,\,(6)$ $0.373$
(4,4,3) $0.347\,\,(5)$ $0.348$ $0.372\,\,(5)$ $0.372$
(4,4,4) $0.344\,\,(8)$ $0.350$ $0.371\,\,(4)$ $0.374$
(Image Deblurring - reconstruction) - For our final example we consider the 2D image reconstruction problem which comes from the software package IRTools [10]. Similarly to the reconstruction of Spectra, we compare the behavior of AIR against MP-IR for various precision choices. The Hubble image for this problem in Figure 3a contains $512\times 512$ pixels which corresponds to a blurring operator of size $512^{2}\times 512^{2}$ making both the application of chop and the storage of the full problem untenable. However, because the point spread function (Figure 3b) is rank-1, the blurring matrix $A$ can be decomposed into a Kronecker product, each part of which is of reasonable size to compute its SVD:
$A=A_{r}\otimes A_{c}=(U_{r}\Sigma_{r}V_{r}^{T})\otimes(U_{c}\Sigma_{c}V_{c}^{T})=(U_{r}\otimes U_{c})(\Sigma_{r}\otimes\Sigma_{c})(V_{r}\otimes V_{c})^{T}.$
With this Kronecker structure, it is possible to compute MP-IR efficiently; see [6] for more implementation details regarding chop and the Kronecker product. To efficiently implement AIR one may form a normal matrix $C$ that admits a spectral factorization by imposing periodic boundary conditions when convolving with the point spread function; see [15] for an overview of 2D image convolution. Because of the astronomical nature of the Hubble image, the choice of periodic or zero boundary conditions is typically experimentally equivalent. Zero boundary conditions were used in the formation of $A$. Figure 3c shows the blurred image with $1\%$ noise.
Figure 3: Image deblurring example: (a) true Hubble image ($512\times 512$ pixels), (b) PSF ($20\times 20$ pixels), (c) blurred and $1\%$ noised image ($512\times 512$ pixels).
Similarly to the numerical results shown for the Spectra reconstruction example, Table 4 provides the RREs at selected iterations and precision combinations for MP-IR as well as AIR as a function of $\alpha^{2}$ and the noise level of the problem. Here we note that we considered different noise levels: $1\%$ and $3\%$ and regularization parameters: 1e0, 1e-1, and 1e-2 compared to the Spectra example, though again for the same reasons as before, we only show results for the two smaller regularization parameters. The same general behaviors observed for the Spectra example were also observed here. In Figure 4 we show the best reconstructions for MP-IR at precision (1,1,1) and (4,4,3) compared to AIR. We note that there are no meaningful visual differences between the images.
Table 4: Image deblurring example: relative errors for MP-IR and AIR methods as a function of the regularization parameter and noise level. The minimum error and its corresponding iteration number in parentheses as well as the final error for the maximum number of iterations allowed (10) is provided. MP-IR errors are provided for various mixed precision combinations.
$\alpha^{2}$ Method (Pr1, Pr2, Pr3) Noise Level
$1\%$ $3\%$
Min. RRE (It.) Fin. RRE Min. RRE (It.) Fin. RRE
1e-1 AIR $0.189\,\,(1)$ $102.7e$ 10 $0.190\,\,(1)$ $108.1e$ 10
MP-IR (1,1,1) $0.189\,\,(1)$ $0.189$ $0.190\,\,(1)$ $0.190$
(2,1,1) $0.189\,\,(3)$ $0.189$ $0.190\,\,(7)$ $0.190$
(2,2,1) $0.189\,\,(3)$ $0.189$ $0.190\,\,(2)$ $0.190$
(3,2,1) $0.189\,\,(3)$ $0.189$ $0.190\,\,(4)$ $0.190$
(3,2,2) $0.189\,\,(4)$ $0.189$ $0.190\,\,(7)$ $0.190$
(3,3,2) $0.189\,\,(3)$ $0.189$ $0.190\,\,(8)$ $0.190$
(4,3,2) $0.189\,\,(3)$ $0.189$ $0.190\,\,(3)$ $0.190$
(4,3,3) $0.189\,\,(10)$ $0.189$ $0.190\,\,(7)$ $0.190$
(4,4,3) $0.189\,\,(3)$ $0.189$ $0.190\,\,(9)$ $0.190$
(4,4,4) $0.189\,\,(4)$ $0.189$ $0.190\,\,(8)$ $0.190$
1e-2 AIR $0.182\,\,(1)$ $192.0e$ 19 $0.243\,\,(1)$ $195.9e$ 19
MP-IR (1,1,1) $0.174\,\,(3)$ $0.174$ $0.184\,\,(4)$ $0.184$
(2,1,1) $0.174\,\,(3)$ $0.174$ $0.184\,\,(10)$ $0.184$
(2,2,1) $0.174\,\,(3)$ $0.174$ $0.184\,\,(3)$ $0.184$
(3,2,1) $0.174\,\,(4)$ $0.174$ $0.184\,\,(6)$ $0.184$
(3,2,2) $0.174\,\,(9)$ $0.174$ $0.184\,\,(6)$ $0.184$
(3,3,2) $0.174\,\,(2)$ $0.174$ $0.184\,\,(10)$ $0.184$
(4,3,2) $0.174\,\,(8)$ $0.174$ $0.184\,\,(7)$ $0.184$
(4,3,3) $0.174\,\,(6)$ $0.174$ $0.184\,\,(5)$ $0.184$
(4,4,3) $0.174\,\,(8)$ $0.174$ $0.184\,\,(7)$ $0.184$
(4,4,4) $0.174\,\,(6)$ $0.174$ $0.184\,\,(7)$ $0.184$
Figure 4: Image deblurring example: (a) MP-IR with precisions $(1,1,1)$ at iteration $7$, (b) AIR at iteration $1$, (c) MP-IR with precisions $(4,4,3)$ at iteration $8$ all using $\alpha^{2}=$1e-2 and at their minimum RRE iteration according to Table 4.
6 Conclusion
In this work, we have investigated IR on the Tikhonov problem in mixed precision. We showed that the iterates computed in mixed precision behave in a filtering manner by deriving methodology to formulate the iterates as a recursive relationship involving the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous terms. Our numerical results showcase that simulated mixed precision IR can give comparable accuracy compared to results computed in double precision and superior results to the described AIR method which is appropriate and efficient to compute with in image deblurring applications.